OpenStudy (anonymous):

How does one find the oblique asymptotes of a rational function? For example in f(x) = (2x^2-3x-1)/(x-2)

OpenStudy (anonymous):

divide it and find the remainder

OpenStudy (anonymous):

i'm looking at here but i can check ur answer < http://answers.yahoo.com/question/index?qid=20090522120106AA53hNZ >

OpenStudy (campbell_st):

use polynomial division and don't worry about the remainder... the quotient will be the asymptote  so y = 2x + 1 is the oblique asymptote.

OpenStudy (anonymous):

I was looking more for why you do that division

OpenStudy (shubhamsrg):

asymptotes are the lines which are tangent to the curve at infinity(+ or -) f(x) = (2x^2-3x-1)/(x-2) let px + q be the asymptote,, so we can see x->infinity(+,-) (f(x) - (px + q)) = 0 => (2x^2-3x-1)/(x-2) - px-q =0 =>(2x+1) + 1/(x-2) -px -q =0 =>1/(x-2) + x(2-p) + 1-q =0 =>now on putting x= infintiy, you can easily see, why p=2 and q=1 or asymp. is 2x + 1.. hope you understand why we did long division..